OpenStudy (anonymous):

The number of eggs,X, laid by the female tawny owl (Strix Aluco) has the probability distribution given in the following table X 2 3 4 P(X=x) 0.1 0.2 0.7 For any egg, the probability that it is hatched is 0.8, independent of all other eggs . Let Y denote the number of hatched eggs in a randomly chosen nest. (a). Obtain the probability distribution of Y (b). Find E(Y) and Var( Y).

OpenStudy (calculusfunctions):

How much were you able to do? Show me your work and then I'll guide you from there.

OpenStudy (tkhunny):

They NEVER lay just one? That's a little odd.

OpenStudy (anonymous):

i was trying to use the method of First value of x and second value of X and combining two sets of results that give the probability distribution of Y, but then i failed to get it

OpenStudy (tkhunny):

Look at these things: (p+q)^2 = ?? (p+q)^3 = ?? (p+q)^4 = ?? See if that leads you anywhere useful.

OpenStudy (unklerhaukus):

\[\sum_{n=1} P_n=1\]

OpenStudy (calculusfunctions):

Do you know the Binomial Distribution?